OpenStudy (anonymous):

in a basketball game, Benito scored 3 points less than twice the number of points Mark scored. Mark scored 8 points. How many points did Benito score?

OpenStudy (anonymous):

B=Benito, M=Mark B=2M-3 M=8 B=2(8)-3 B=16-3 B=13